OpenStudy (anonymous):

the press box at a baseball park is 33.0 ft above the ground. A reporter in the press box looks at an angle of 19.0 degrees below the horizontal to see second base. what is the horizontal distance form the press box to second base? This is what i did.... 90+19.0=109 180-109=71 (sin 90x)(sin 19/33) =.4867547088 I THINK I'AM WRONG. CAN SOMEONE HELP ME! PLEASE. THANKS

OpenStudy (ghazi):

why you added 90 to 19

OpenStudy (ghazi):

33/x = cos 19

OpenStudy (ghazi):

x= 33/ cos 19

OpenStudy (ghazi):

y/x = sin 19...........y= x*sin 19= 35.10*0.32=11.2 ft from the first base

OpenStudy (ghazi):



OpenStudy (anonymous):

yeah the answer actually is 95.838!! thanks.

OpenStudy (ghazi):

:)